OpenStudy (noseboy908):

Pre-Calculus: Find the missing term of the arithmetic sequence: 19,_, _,-71

OpenStudy (anonymous):

The missing TERMS or term

OpenStudy (noseboy908):

Oh, sorry. It should have read "terms".

OpenStudy (anonymous):

Arithmetic sequence \[A_n= A_1 + (n-1)d\]

OpenStudy (anonymous):

First you would have to find the difference which i believe can be done by using -71

OpenStudy (anonymous):

Since -71 is the 4th term then you can solve for d in the formula

OpenStudy (anonymous):

\[-71_{4th} = 19 + (4th-1)d\]

OpenStudy (noseboy908):

Alright, I'll try it and find d.

OpenStudy (noseboy908):

-71=19+(4-1)d -71=3d+19 -90=3d d=-30

OpenStudy (noseboy908):

Thanks!

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Now you can solve for the missing terms with the difference known. Great Job!